Filer Manufacturing has 8 million shares of common stock outstanding. The current share price is $74, and the book value per share is $5. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $80 million and a coupon rate of 9 percent and sells for 95 percent of par. The second issue has a face value of $60 million and a coupon rate of 10 percent and sells for 108 percent of par. The first issue matures in 24 years, the second in 8 years.


a.

What are the company’s capital structure weights on a book value basis? (Do not round intermediate calculations. Round your answers to 4 decimal places (e.g., 32.1616).)


  Equity/Value    
  Debt/Value    


b.

What are the company’s capital structure weights on a market value basis? (Do not round intermediate calculations. Round your answers to 4 decimal places (e.g., 32.1616).)


    Equity/Value    
    Debt/Value    


c.

Which are more relevant, the book or market value weights?

  • Market value

  • Book value
Answer #1

a

BV of equity=Price of equity*number of shares outstanding
BV of equity=5*8000000
=40000000
BV of Bond1=Par value*bonds outstanding*%age of par
BV of Bond1=1000*80000*1
=80000000
BV of Bond2=Par value*bonds outstanding*%age of par
BV of Bond2=1000*60000*1
=60000000
BV of firm = BV of Equity + BV of Bond1+ BV of Bond 2
=40000000+80000000+60000000
=180000000
Weight of equity = BV of Equity/BV of firm
Weight of equity = 40000000/180000000
W(E)=0.2222
Weight of debt = BV of Bond/BV of firm
Weight of debt = 140000000/180000000
W(D)=0.7778

b

MV of equity=Price of equity*number of shares outstanding
MV of equity=74*8000000
=592000000
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=1000*80000*0.95
=76000000
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=1000*60000*1.08
=64800000
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2
=592000000+76000000+64800000
=732800000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 592000000/732800000
W(E)=0.8079
Weight of debt = MV of Bond/MV of firm
Weight of debt = 140800000/732800000
W(D)=0.1921

c